he length of a rectangle is 2 more than twice its width. The perimeter is 36 feet
Answer by blwinbbbles(106)

Write down what you know: w = width l=length
you know that its perimeter is 36 so:
2w + 2l = 36
and you know l = 2 + 2w so you can use substitution since you know that l is equal to 2+2w
2w + 2(2 + 2w) = 36
2w + 4 + 4w = 36
6w + 4 = 36 subtract both sides by 4
6w = 32 divide by 6
w = 5.33
now sub 5.33 in for w
2 +2(5.33) = l
2 + 10.66 = 12.66
so length is 12.66 and width is 5.33
Hope this helps